# Sufficient digits and density estimation: A Bayesian nonparametric approach using generalized finite Pólya trees

### v1

@@ 7.2 Posterior results when using GFPT1 and GFPT2 priors
This section investigates through simulation studies if there is any gain in extending GFPT1 models to GFPT2 models. For this we consider six different data generating processes, namely DG1–DG3 as in Section 7.1 and
@@
$Z_{1},\ldots,Z_{m}$ $\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\frac{1}{2}\mathrm{B}(2,15)+\frac{1}{2}\mathrm{B}(15,2).$ (DG6)
@@
We consider a GFPT1 prior where $\alpha_{0}=0.1$ (as suggested at the end of Section 7.1) and a GFPT2 prior where $\alpha_{x_{1:n}}=\alpha_{0}n^{2}$ with $\alpha_{0}\in\{0.1,2.0\}$ (in Section 7.1, these values of $\alpha_{0}$ provided the best GFPT1 posterior estimates). As far as the prior for $R$ is concerned (cf. Defintion 4.5), we assume $\beta_{x_{1:n}}=\beta_{0}$ for all $n\in\mathbb{N}$ and all $x_{1:n}\in\{0,1\}^{n}$, so that the prior mean for the NBP $I$ is given by the standard dyadic partitions employed by the GFPT1 model. We let $\beta_{0}\in\{0.5,2.0,5.0\}$, which allowed us to understand the sensitivity of GFPT2 priors: When $\beta_{0}=0$, the $\mathrm{BS}(\beta)$ prior for $R$ assigns significant mass to values of $R_{x_{1:n}}$ near to 00 or $1$, corresponding to almost empty intervals. Instead when $\beta_{0}\in\{2.0,5.0\}$, the $\mathrm{BS}(\beta)$ prior for $R$ causes a random NBP which is more and more concentrated around the standard dyadic partitions. We also tried increasing the value of of $\beta_{0}$ when increasing the level $n$ of the partitions, but this led to worse mixing of the MCMC algorithm for posterior simulations (see Remark 13; we used 10,000 iterations and discarding the first 1,000 as burn-in). In our examples, running this MCMC algorithm on a standard laptop takes only around five seconds.
@@
Figure 5: Total variation distance (TV) between true and estimated densities (first two rows of plots) and posterior expectation of $N$ (last two rows). Each boxplot summarizes the results from 100 independent replicates of the indicated data generating model. The $\beta_{0}$ parameter of the GPFT2 model varies across the $x$-axis of each plot.
@@
For each data generating process, we simulated 1,000 observations and repeated each analysis on 100 independently simulated datasets. Figure 5 shows for the different data generating processes and parameter values, the total variation distance between true and estimated densities as well as the posterior mean of $N$. We observer the following. First, the parameter $\beta_{0}$ does not seem to play a significant role in posterior inference. Second, when using GFPT2 priors, as in the case of GFPT1 priors, higher values of $\alpha_{0}$ correspond to smaller values of $N$ a posteriori. In some scenarios (e.g., DG1 and DG4) this yields also a more accurate estimation of the density because the density needs only a coarse partition to be well approximated. In other scenarios (e.g., DG3 and DG6), this results in a much poorer density estimate. In both such cases, the data generating density entails that the number of sufficient digits is infinite, and a finer partition results in better density estimates. Focusing specifically on DG4 and DG5, observe that those data generating densities are piecewise continuous on intervals that do not agree with the standard diadic partitions employed for the GFPT1 prior. As a result, under the GFPT1 prior, a posterior $N$ tends to be large in those scenarios, while especially when $\alpha_{0}=2$ the GFPT2 model correctly captures the partitions needed to approximate those densities with a small number of intervals, resulting in smaller $N$ a posteriori. In particular, for DG4, using the GFPT2 prior with $\alpha_{0}=2$ achieves almost a perfect density estimate.
@@ 7.3 Fitted models for real datasets spanning multiple orders of magnitude
Data for the first four dataset can be found at https://github.com/jasonlong/benfords-law, while data for the income dataset is publicly available from https://www.census.gov/data.html. Table 1 reports the empirical range (maximum and minimum value of the data) in the different datasets. The datasets in Table 1 span four, three, eight, six, and seven orders of magnitude, respectively, making them particularly well suited to test the adequacy of the base-2 and base-10 multiscale Benford Pólya tree priors.
@@
Table 1: Empirical range of the observations in the different datasets analyzed in Section 7.3.
@@
We compare posterior results when using five different priors, namely $\mathrm{MBPT}_{2}$ and $\mathrm{MBPT}_{10}$ priors, GFPT1 and GFPT2 priors fitted on max-scaled data (i.e., for each dataset, we divide all observations by the maximum value), and a Dirichlet process mixture (DPM) of Gaussian densities [11, 14, 28, 32], which is the de-facto standard model for Bayesian density estimation. The parameters of the GFPT1 and GFPT2 priors are selected following Section 4.3 and the insights developed in Section 7.1. Specifically, $\alpha_{x_{1:n}}=0.1n^{2}$ for both priors, and $\beta_{x_{1:n}}=02$ for the CFPT2 prior. For the prior on $N$, we proceed as follows. Let $n_{\max}$ be the maximum number of base-10 digits recorded in a given dataset. Then, for the $\mathrm{MBPT}_{10}$ prior, assume that $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,n_{\max}\}$, where the Poisson distribution has mean $n_{\max}/2$. For the $\mathrm{MBPT}_{2}$, GFPT1, and the GFPT2 priors, assume that $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,\lceil\log_{2}(10^{n_{\max}})\rceil\}$, where the Poisson distribution has mean $\log_{2}(10^{n_{\max}})/2$. This ensures that under all five priors of $N$, the number of possible values of the sufficient digits is approximately the same. Further, for the $\mathrm{MBPT}_{2}$ and $\mathrm{MBPT}_{10}$ priors, we fix $T$ as the span of the orders of magnitudes of the data, set $\omega_{k}\propto 1$ for all $k$, and let $c_{n}$ from (18) be given by $c_{n}=c_{0}n^{2}$ for $n\in\mathbb{N}$, where $c_{0}=0.1$ in case of $\mathrm{MBPT}_{10}$ and $c_{0}=2.0$ in case of $\mathrm{MBPT}_{2}$. This choice of $c_{0}$ entails that the variance of the random PDF in (6) is approximately equal under both MBPT priors. Finally, for the DPM prior, we use the implementation and follow the default prior elicitation strategy in the BayesMix library [3].
@@
The posterior inferences using the different priors are compared in term of the widely applicable information criterion (WAIC), see [37], which is a consistent estimator of the out-of-sample error [36]. Since the absolute value of the WAIC is irrelevant, we fix the DPM as a baseline prior and report in Table 2 the relative improvement of the WAIC for any model over the DPM (larger values of the relative improvement correspond to better model performance). For datasets eurodist, GDP, and census, all priors seem to perform similarly. For the datasets Twitter and income datasets, the multiscale Benford Pólya tree priors clearly outperforms all other priors, where in particular, the $\mathrm{MBPT}_{10}$ prior shows a remarkable improvement of rougly 19% and 25% over the DPM prior, respectively. Moreover, in all scenarios, the WAIC when using the GFPT2 prior slightly improves as compared to the GFPT1 prior.
@@
[…]
@@ 8 Concluding remarks and open problems
A natural extensions is to consider multivariate data, either by defining a product nested binary partition on the unit cube, or by relying on coordinate values given by space-filling curves such as Morton’s or Hilbert’s curves. Another idea which we leave for future research is the slight modification of the GFPT2 model obtained by the mixture distribution
$R_{x_{1:n}}\,|\,R_{x_{1:n-1}}\sim\begin{cases}\pi\delta_{\{1\}}+(1-\pi)B(\beta_{x_{1:n-1},0},\beta_{x_{1:n-1},1})&\text{if }R_{x_{1:n-1}}<1\\ \delta_{\{1\}}&\text{if }R_{x_{1:n-1}}=1\end{cases}$
where $10<\pi<$ 1 is a parameter. In this way, we introduce an “optional” stopping of the partitioning in some regions of the domain, similarly to what is done in [38].
@@
Ongoing work by us extends the setting in this paper, utilizing the numeric system induced by a continued fraction representation. That is, $X_{1},X_{2},...\in\mathbb{N}$ are the random digits of $X$ such that $X=1/(X_{1}+(1/X_{2}+...))$. Thus, the digits take value in an unbounded space, which complicates inference, yet promises sharper inference because continued fractions give the best rational approximations of real numbers.
@@ Acknowledgements
Supported by The Danish Council for Independent Research — Natural Sciences, grant DFF – 10.46540/2032-00005B.
@@ Appendix A Proofs for Section 3
The existence and coupling of $N$ with $X$ such that (3) is satisfied follows from [29].
Let the situation be as in the second part of the theorem. Clearly, $g$ is a PDF on $H$. Recall that a set $U\subseteq H$ is open (with respect to $H$) if it is the intersection of $H$ with an open subset of $\mathbb{R}$. The following facts are well-known, see e.g. [10].
@@
Let $n\in\mathbb{N}_{0}$, $y^{(n)}=0$ if $n=0$, and $y^{(n)}=(y_{x_{1:j}}\,|\,x_{1:j}\in\{0,1\}^{j},j=1,\ldots,n)$ if $n>0$ where $0\leq y_{x_{1:j-1},0}\leq 1$ and $y_{x_{1:j-1},1}=1-y_{x_{1:j-1},0}$. Then we can identity $y^{(n)}$ by $y^{(n)}_{0}=(y_{x_{1:j-1},0}\,|\,x_{1:j-1}\in\{0,1\}^{j-1},1\leq j\leq n)$. Definitions 4.1 and 4.3 give that the distribution of $Y^{(N)}$ (or more precisely $Y^{(N)}_{0}$) is absolutely continuous with respect to $\mu$, with density
@@
The conditional distribution of $N$ given $Z_{1:m}=z_{1:m}$ has a PMF $p(n\,|\,z_{1:m})$ with $p(0\,|\,z_{1:m})\propto p(0)$ and where for every $n\in\mathbb{N}$, (B1) gives
@@
which reduces to (20). It also follows from (B1) that the conditional distribution of $Y^{(N)}$ given $(N,Z_{1:m})=(n,z_{1:m})$ is absolutely continuous with respect to $\mu_{n}$, with a density $p(y\,|\,n,z_{1:m})$ for $y\in A_{n}$ so that if $n>0$,
@@
To establish (C2) we use the following facts, considering any $n\in\mathcal{N}$ such that $n+1\in\mathcal{N}$. By the law of large numbers,
@@
×$\hskip 56.9055pt\times\frac{1}{\mathrm{b}(\alpha_{x_{1:n},0}+N_{x_{1:n},0},\alpha_{x_{1:n},1}+N_{x_{1:n},1})}.$ (C4)
@@
$\frac{p(n\,|\,Z_{1:m})}{p(n+1\,|\,Z_{1:m})}\asymp{\frac{p_{n}}{p_{n+1}}}\times$
$\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{(\ell_{x_{1:n},0}+\ell_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{-1/2}}}$ (C5)
thanks to (C3), (C4), the condition on the $\alpha_{x_{1:n}}$’s, and Stirling’s approximation of the beta function: $\mathrm{b}(x,y)=\sqrt{2\pi}\frac{x^{x-1/2}y^{y-1/2}}{(x+y)^{x+y-1/2}}\left(1+O(1/x)+O(1/y)\right)$.
Now, consider any integer $n\in\mathcal{N}$ with $n>n^{*}$. By Theorem 3.2, for any $y\in I_{x_{1:n}}$ we have that $f^{*}(y)$ is constant and $\int_{I_{x_{1:n}}}f^{*}(v)\,\mathrm{d}v=f^{*}(y)\ell_{x_{1:n}}$. Hence, for $k=0,1$ and $i_{x_{1:n}}=1$, it follows from (C3) that
@@
×$\times\left(\frac{N_{x_{1:n},0}+N_{x_{1:n},1}}{N_{x_{1:n},0}}\right)^{N_{x_{1:n},0}-1/2}\left(\frac{N_{x_{1:n},0}+N_{x_{1:n},1}}{N_{x_{1:n},1}}\right)^{N_{x_{1:n},1}-1/2}$
×$\times[(N_{x_{1:n},0}+N_{x_{1:n},1})^{1/2}/{\sqrt{2\pi}}]$
$\asymp{\frac{p_{n}}{p_{n+1}}}\prod_{x_{1:n}\in\{0,1\}^{n}:\,i_{x_{1:n}}=1}(N_{x_{1:n}}/{(2\pi)})^{1/2}.$ (C7)
By (30), the term $p_{n}/p_{n+1}$ is strictly positive and bounded, so (C7) gives
@@
Let $\tilde{n}\in\mathcal{N}$ with $\tilde{n}>n^{*}$. By (30), $\{\tilde{n},\tilde{n}-1,\ldots,n^{*}\}\subseteq\mathcal{N}$. Hence, by (C8),
@@
whereby (C2) is verified.
Consider instead any $n\in\mathcal{N}$ with $n<n^{*}$. Define $d_{x_{1:n}}=0$ if $f^{*}$ is constant over $I_{x_{1:n}}$ and $d_{x_{1:n}}=1$ otherwise. Then (C5) writes

### v2

@@ 6.2 Posterior results when using GFPT1 and GFPT2 priors
This section investigates through simulation studies if there is any gain in extending GFPT1 models to GFPT2 models. Moreover, we also compare our models with the Optional Pólya tree (OPT) model of Wong and Ma (2010) and the APT model of Ma (2017), implemented in the R package PTT. For this we consider six different data generating processes, namely DG1–DG3 as in Section 6.1 and
@@
$Z_{1},\ldots,Z_{m}$ $\stackrel{{\scriptstyle\mbox{\scriptsize iid}}}{{\sim}}\frac{1}{2}\mathrm{Beta}(2,15)+\frac{1}{2}\mathrm{Beta}(15,2).$ (DG6)
@@
We consider a GFPT1 prior where $\alpha_{0}=0.1$ (as suggested at the end of Section 6.1) and a GFPT2 prior where $\alpha_{x_{1:n}}=\alpha_{0}n^{2}$ with $\alpha_{0}\in\{0.1,2.0\}$ (in Section 6.1, these values of $\alpha_{0}$ provided the best GFPT1 posterior estimates). As far as the prior for $R$ is concerned (cf. Defintion 4.6), we assume $\beta_{x_{1:n}}=\beta_{0}$ for all $n\in\mathbb{N}$ and all $x_{1:n}\in\{0,1\}^{n}$, so that the prior mean for the NBP $I$ is given by the standard dyadic partitions employed by the GFPT1 model. We let $\beta_{0}\in\{0.5,2.0,5.0\}$, which allowed us to understand the sensitivity of GFPT2 priors: When $\beta_{0}=0.5$, the $\mathrm{BS}(\beta)$ prior for $R$ assigns significant mass to values of $R_{x_{1:n}}$ near to 0 or $1$, corresponding to almost empty intervals. Instead when $\beta_{0}\in\{2.0,5.0\}$, the $\mathrm{BS}(\beta)$ prior for $R$ causes a random NBP which is more and more concentrated around the standard dyadic partitions. We also tried increasing the value of of $\beta_{0}$ when increasing the level $n$ of the partitions, but this led to worse mixing of the MCMC algorithm for posterior simulations (see Remark 14; we used 10,000 iterations and discarding the first 1,000 as burn-in). In our examples, running this MCMC algorithm on a standard laptop takes only around five seconds. The prior specification for the OPT and APT models follows the defaults of the PTT package.
@@
Figure 7: Total variation distance (TV) between true and estimated densities. Each boxplot summarizes the results from 100 independent replicates of the indicated data generating model.
Figure 8: Posterior expectation of $N$ for the simulated datasets in Section 6.2. Each boxplot summarizes the results from 100 independent replicates of the indicated data generating model.
@@
For each data generating process, we simulated 1,000 observations and repeated each analysis on 100 independently simulated datasets. Figure 7 shows for the different data generating processes and parameter values, the total variation distance between true and estimated densities obtained with the different models. Moreover, Figure 8 shows the posterior mean of $N$ under the GFPT1 and GPFT2 models. We observer the following. First, the parameter $\beta_{0}$ does not seem to play a significant role in posterior inference. Second, when using GFPT2 priors, as in the case of GFPT1 priors, higher values of $\alpha_{0}$ correspond to smaller values of $N$ a posteriori. In some scenarios (e.g., DG1 and DG4) this yields also a more accurate estimation of the density because the density needs only a coarse partition to be well approximated. In other scenarios (e.g., DG3 and DG6), this results in a much poorer density estimate. In both such cases, the data generating density entails that the number of sufficient digits is infinite, and a finer partition results in better density estimates. Focusing specifically on DG4 and DG5, observe that those data generating densities are piecewise continuous on intervals that do not agree with the standard diadic partitions employed for the GFPT1 prior. As a result, under the GFPT1 prior, a posterior $N$ tends to be large in those scenarios, while especially when $\alpha_{0}=2$ the GFPT2 model correctly captures the partitions needed to approximate those densities with a small number of intervals, resulting in smaller $N$ a posteriori. In particular, for DG4, using the GFPT2 prior with $\alpha_{0}=2$ achieves almost a perfect density estimate. Finally, the OPT and APT models yield comparable performance to the GFPT2 model in all settings except when data are generated from DG4 and DG5, in which case the GFPT2 models achieve superior density estimates.
@@ 6.3 Fitted models for real datasets spanning multiple orders of magnitude
Data for the first four dataset can be found at https://github.com/jasonlong/benfords-law, while data for the income dataset is publicly available from https://www.census.gov/data.html. Table 1 reports, for each dataset, the empirical range (maximum and minimum value of the data) and two diagnostics computed from the log-mantissa $U=\log_{10}(Z)-\lfloor\log_{10}(Z)\rfloor$: The Wasserstein-1 distance $W_{1}$ between the empirical distribution of $U$ and the $\mathrm{Unif}[0,1)$ law, and the Kolmogorov-Smirnov distance $d_{\mathrm{KS}}$ to $\mathrm{Unif}[0,1)$. Under the Newcomb-Benford law, the distribution is scale invariant, i.e., the log-mantissa $U$ is $\mathrm{Unif}[0,1)$-distributed, hence, small values of $W_{1}$ and $d_{\mathrm{KS}}$ indicate that our MBPT model is well supported by the data. The datasets in Table 1 span four, three, eight, six, and seven orders of magnitude, respectively, making them particularly well suited to test the adequacy of the base-2 and base-10 multiscale Benford Pólya tree priors. In addition, the $W_{1}$ and $d_{\mathrm{KS}}$ statistics quantify how close each dataset is to the Benford benchmark: eurodist has the largest deviations (0.18 and 0.07), signalling weak scale-invariance; by contrast, census is extremely close to uniform (0.001 and 0.003), with income, GDP, and Twitter also showing relatively small deviations. This pattern anticipates our empirical findings, whereby the MBPT model delivers the best fit where the data spans a large number of orders of magnitude and $W_{1}$ and $d_{\mathrm{KS}}$ are close to zero.
@@
Table 1: Empirical range of the observations and $W_{1}$ and $d_{\mathrm{KS}}$ diagnostics in the different datasets analyzed in Section 6.3.
@@
We compare posterior results when using five different priors, namely $\mathrm{MBPT}_{2}$ and $\mathrm{MBPT}_{10}$ priors, GFPT1 and GFPT2 priors fitted on max-scaled data (i.e., for each dataset, we divide all observations by the maximum value), and a Dirichlet process mixture (DPM) of Gaussian densities (Escobar and West, 1995; Ghoshal and Vaart, 2017; MacEachern, 1994; Neal, 2000), which is the de-facto standard model for Bayesian density estimation. The parameters of the GFPT1 and GFPT2 priors are selected following Section 4.3 and the insights developed in Section 6.1. Specifically, $\alpha_{x_{1:n}}=0.1n^{2}$ for both priors, and $\beta_{x_{1:n}}=02$ for the CFPT2 prior. For the prior on $N$, we proceed as follows. Let $n_{\max}$ be the maximum number of base-10 digits recorded in a given dataset. Then, for the $\mathrm{MBPT}_{10}$ prior, assume that $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,n_{\max}\}$, where the Poisson distribution has mean $n_{\max}/2$. For the $\mathrm{MBPT}_{2}$, GFPT1, and the GFPT2 priors, assume that $N$ follows the truncation of a Poisson distribution to $\{0,1,\ldots,\lceil\log_{2}(10^{n_{\max}})\rceil\}$, where the Poisson distribution has mean $\log_{2}(10^{n_{\max}})/2$. This ensures that under all five priors of $N$, the number of possible values of the sufficient digits is approximately the same. Further, for the $\mathrm{MBPT}_{2}$ and $\mathrm{MBPT}_{10}$ priors, we fix $T$ as the span of the orders of magnitudes of the data, set $\omega_{k}\propto 1$ for all $k$, and let $c_{n}$ from (18) be given by $c_{n}=c_{0}n^{2}$ for $n\in\mathbb{N}$, where $c_{0}=0.1$ in case of $\mathrm{MBPT}_{10}$ and $c_{0}=2.0$ in case of $\mathrm{MBPT}_{2}$. This choice of $c_{0}$ entails that the variance of the random PDF in (6) is approximately equal under both MBPT priors. Finally, for the DPM prior, we use the implementation and follow the default prior elicitation strategy in the BayesMix library (Beraha et al., 2022).
@@
The posterior inferences using the different priors are compared in term of the widely applicable information criterion (WAIC), see Watanabe (2013), which is a consistent estimator of the out-of-sample error (Vehtari et al., 2017). Since the absolute value of the WAIC is irrelevant, we fix the DPM as a baseline prior and report in Table 2 the relative improvement of the WAIC for any model over the DPM (larger values of the relative improvement correspond to better model performance, while negative values indicate that the DPM achieves a better fit to the data). For datasets eurodist, GDP, and census, all priors seem to perform similarly. For the datasets Twitter and income datasets, the multiscale Benford Pólya tree priors clearly outperforms all other priors, where in particular, the $\mathrm{MBPT}_{10}$ prior shows a remarkable improvement of rougly 19% and 25% over the DPM prior, respectively. These results confirm the intuition obtained by looking at the diagnostics based on the log-mantissa and the span of the datasets discussed above. Moreover, in all scenarios, the WAIC when using the GFPT2 prior slightly improves as compared to the GFPT1 prior. Although not reported here, we have also tried fitting a variant of the MBPT models where the parameters $\alpha$ are not specified in accordance to the Newcomb-Benford law but according to the standard choice of (16). We noticed substantial deterioration in the WAICs across all datasets, and in particular for Twitter, for which this variant performs worse than the DPM baseline, and income, for which this variant performs on par to the baseline.
We now try to explain the reasons behind the large improvements of the $\mathrm{MBPT}_{10}$ model Twitter, income and census data. For Twitter and income, we believe that the main driver is human-driven bunching around thresholds: For instance, Twitter users are incentivized to reach a given number of followers either for monetization (e.g., power of ten followers) or for psychological reasons. Similarly, salary bands are usually set just above psychological or contractual cut-offs. For the census data instead, there might be population counts clustering near administrative thresholds for municipal classification. Such repeating motifs are naturally expressed as stable digit patterns within each order of magnitude, which is what our MBPT models enforce.
@@
Finally, we comment on the number of sufficient digits needed in the different datasets according to the GFPT and MBPT models. Table 3 reports the (estimated) a posteriori expectation of $N$ in the different models. Note that $N$ is typically small a posteriori, despite $n_{\max}$ being equal to six, four, seventeen, eight, and six in the different datasets respectively. Recall that the prior for $N$ is trucated at $n_{\max}$ for the $\mathrm{MBPT}_{10}$ prior and to $\lceil\log_{2}(10^{n_{\max}})\rceil$ for the other priors. Hence, the sufficient number of digits are effectively always less than $n_{\max}$. Except for the Twitter and income datasets, we observe that $N$ tends to be smaller a posteriori under the $\mathrm{MBPT}_{2}$ prior than under the GFPT1 and GFPT2 priors. This entails that the max-scaling of the data disaligns the data from the nested binary partitions. Indeed, using the GFPT2 prior, the posterior adaptively learns the NBP, leading to a smaller $N$ than under the GFPT1 prior that works with the fixed NBP. For all the different datasets, a posteriori $N$ tends to be smaller under the $\mathrm{MBPT}_{10}$ prior, however, the possible values taken by $X_{1:N}$ is $10^{N}$ under the $\mathrm{MBPT}_{10}$ prior and (approximately) $2^{N}$ under the other models. We report the corresponding (estimated) posterior expectations in Table 3. When using the $\mathrm{MBPT}_{10}$ prior a larger number of possible digits is required a posteriori than if using another of the priors, but compared with the $\mathrm{MBPT}_{2}$ prior, the difference may not be extreme, since the scale is coarser for $10^{N}$ than for $2^{N}$. On the other hand, there are several situations in which the $\mathrm{GPFT2}$ prior causes a posteriori a much lower number of possible digits than if using another of the priors, thanks to the added flexibility of learning the NBP.
@@ 7 Concluding remarks and open problems
A natural extensions is to consider multivariate data. Two simple approaches would be to define a product nested binary partition on the unit cube and model each dimension independently with a GFPT1 or GFPT2 prior, or to rely on coordinate values given by space-filling curves such as Morton’s or Hilbert’s curves. However, we anticipate that even in these simple extensions, more efficient algorithms for posterior inference would need to be devised in order to scale to high-dimensional settings, perhaps borrowing ideas from Awaya and Ma (2024). On the other hand, more complex models would be needed to infer the dependence across dimensions. We leave this problems as interesting ideas for future research. Moreover, we also plan to investigate a slight modification of the GFPT2 model obtained by the mixture distribution
[…]
where $0<\pi<1$ is a parameter. In this way, we introduce an “optional” stopping of the partitioning in some regions of the domain, similarly to what is done in Wong and Ma (2010).
@@
Ongoing work by us extends the setting in this paper, utilizing the numeric system induced by a continued fraction representation. That is, $X_{1},X_{2},$$\in\mathbb{N}$ are the random digits of $X$ such that $)X=1/(X_{1}+(1/X_{2}+$...)). Thus, the digits take value in an unbounded space, which complicates inference, yet promises sharper inference because continued fractions give the best rational approximations of real numbers.
@@ Acknowledgements
Mario Beraha gratefully acknowledges support from the Italian Ministry of Education, University and Research (MUR), “Dipartimenti di Eccellenza” grant 2023-2027. Jesper Møller is upported by The Danish Council for Independent Research — Natural Sciences, grant DFF – 10.46540/2032-00005B.
@@
The existence and coupling of $N$ with $X$ such that (3) is satisfied follows from Møller (2024).
Let the situation be as in the second part of the theorem. Clearly, $g$ is a PDF on $H$. Recall that a set $U\subseteq H$ is open (with respect to $H$) if it is the intersection of $H$ with an open subset of $\mathbb{R}$. The following facts are well-known, see e.g. Edwards (1995).
@@
Appendix A Proofs for Section 4.1
Proof of Proposition 4.5.
By the total law of expectation and since $N$ and $Y$ are independent,
$\mathrm{E}[P(I_{x_{1:d}})\,|\,Y]=\sum_{n=0}^{\infty}p_{n}\mathrm{E}[P(I_{x_{1:d}}\,|\,Y,N=n)]$
where
$\mathrm{E}[P(I_{x_{1:d}}\,|\,Y,N=n)]=\begin{cases}\prod_{j=1}^{d}Y_{x_{1:j}}&\text{if }n\geq d,\\ \frac{\ell_{x_{1:d}}}{\ell_{x_{1:n}}}\prod_{j=1}^{n}Y_{x_{1:j}}&\text{if }n<d.\end{cases}$
Taking expectations with respect to $Y$ leads to the result. ∎
@@
[…]
@@
The conditional distribution of $N$ given $Z_{1:m}=z_{1:m}$ has a PMF $p(n\,|\,z_{1:m})$ with $p(0\,|\,z_{1:m})\propto p(0)$ and where for every $n\in\mathbb{N}$, (B1) in this supplementary material gives
@@
which reduces to (20). It also follows from (B1) in this supplementary material that the conditional distribution of $Y^{(N)}$ given $(N,Z_{1:m})=(n,z_{1:m})$ is absolutely continuous with respect to $\mu_{n}$, with a density $p(y\,|\,n,z_{1:m})$ for $y\in A_{n}$ so that if $n>0$,
@@
To establish (C2) in this supplementary material we use the following facts, considering any $n\in\mathcal{N}$ such that $n+1\in\mathcal{N}$. By the law of large numbers,
@@
$\hskip 56.9055pt\times\frac{1}{\mathrm{b}(\alpha_{x_{1:n},0}+N_{x_{1:n},0},\alpha_{x_{1:n},1}+N_{x_{1:n},1})}.$ (C4)
@@
$\frac{p(n\,|\,Z_{1:m})}{p(n+1\,|\,Z_{1:m})}\asymp{\color[rgb]{0,0,0}\frac{p_{n}}{p_{n+1}}}\times$
$\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{(\ell_{x_{1:n},0}+\ell_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\color[rgb]{0,0,0}\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{\color[rgb]{0,0,0}-1/2}}}$ (C5)
thanks to equations (C3) and (C4) in this supplementary material, the condition on the $\alpha_{x_{1:n}}$’s, and Stirling’s approximation of the beta function:
$\mathrm{b}(x,y)=\sqrt{2\pi}\frac{x^{x-1/2}y^{y-1/2}}{(x+y)^{x+y-1/2}}\left(1+O(1/x)+O(1/y)\right).$
Now, consider any integer $n\in\mathcal{N}$ with $n>n^{*}$. By Theorem 3.2, for any $y\in I_{x_{1:n}}$ we have that $f^{*}(y)$ is constant and $\int_{I_{x_{1:n}}}f^{*}(v)\,\mathrm{d}v=f^{*}(y)\ell_{x_{1:n}}$. Hence, for $k=0,1$ and $i_{x_{1:n}}=1$, it follows from (C3) in this supplementary material that
@@
$\times\left(\frac{N_{x_{1:n},0}+N_{x_{1:n},1}}{N_{x_{1:n},0}}\right)^{N_{x_{1:n},0}-1/2}\left(\frac{N_{x_{1:n},0}+N_{x_{1:n},1}}{N_{x_{1:n},1}}\right)^{N_{x_{1:n},1}-1/2}$
$\times[(N_{x_{1:n},0}+N_{x_{1:n},1})^{1/2}/{\color[rgb]{0,0,0}\sqrt{2\pi}}]$
$\asymp{\color[rgb]{0,0,0}\frac{p_{n}}{p_{n+1}}}\prod_{x_{1:n}\in\{0,1\}^{n}:\,i_{x_{1:n}}=1}(N_{x_{1:n}}/{(2\pi)})^{1/2}.$ (C7)
By (30), the term $p_{n}/p_{n+1}$ is strictly positive and bounded, so (C7) in this supplementary material gives
@@
Let $\tilde{n}\in\mathcal{N}$ with $\tilde{n}>n^{*}$. By (30), $\{\tilde{n},\tilde{n}-1,\ldots,n^{*}\}\subseteq\mathcal{N}$. Hence, by (C8) in this supplementary material,
@@
whereby (C2) in this supplementary material is verified.
Consider instead any $n\in\mathcal{N}$ with $n<n^{*}$. Define $d_{x_{1:n}}=0$ if $f^{*}$ is constant over $I_{x_{1:n}}$ and $d_{x_{1:n}}=1$ otherwise. Then (C5) in this supplementary material writes